### v2

# Uncertainty Quantification for Retrieval-Augmented Reasoning

Abstract.
Retrieval-augmented reasoning (RAR) is a recent evolution of retrieval-augmented generation (RAG) that employs multiple reasoning steps for retrieval and generation. While effective for some complex queries, RAR remains vulnerable to errors and misleading outputs. Uncertainty quantification (UQ) offers methods to estimate the confidence of systems’ outputs. These methods, however, often handle simple queries with no retrieval or single-step retrieval, without properly handling RAR setup. Accurate estimation of UQ for RAR requires accounting for all sources of uncertainty, including those arising from retrieval and generation. In this paper, we account for all these sources and introduce Retrieval-Augmented Reasoning Consistency ($\text{R}^{2}$C)–a novel UQ method for RAR. The core idea of $\text{R}^{2}$C is to perturb the multi-step reasoning process by applying various actions to reasoning steps. These perturbations alter the retriever’s input, which shifts its output and consequently modifies the generator’s input at the next step. Through this iterative feedback loop, the retriever and generator continuously reshape one another’s inputs, enabling us to capture uncertainty arising from both components. Experiments on five popular RAR systems across diverse QA datasets show that $\text{R}^{2}$C improves AUROC by over 5% on average compared to the state-of-the-art UQ baselines. Extrinsic evaluations using $\text{R}^{2}$C as an external signal further confirm its effectiveness for two downstream tasks: in Abstention, it achieves ~5% gains in both F1Abstain and AccAbstain; in Model Selection, it improves the exact match by ~7% over single models and ~3% over selection methods.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Figure 1. $\text{R}^{2}$C overview. Given a user query, the agent (LLM) first generates the most-likely reasoning path leading to the most-likely response (left, yellow). To estimate uncertainty, $\text{R}^{2}$C creates multiple perturbed generations by randomly altering states in the reasoning path (middle, gray). The confidence score is then derived via majority voting. $\text{R}^{2}$C significantly outperforms established UQ methods and achieves significant improvements on two downstream tasks: abstention and model selection.
Retrieval-augmented generation (RAG) is widely used for knowledge-intensive tasks, but remains limited in addressing complex multi-step reasoning (Jin et al., 2025; Soudani et al., 2024b; Salemi and Zamani, 2024; Soudani et al., 2024c). Recent work has explored combining RAG with reasoning, where LLMs are prompted or trained to use search engines as tools during their reasoning process; a paradigm referred to as retrieval-augmented reasoning (RAR) (Sun et al., 2025; Tran et al., 2025; Islam et al., 2024; Jin et al., 2025). However, RAR models are still prone to producing incorrect responses, due to issues such as retrieving irrelevant documents in early steps, misinterpreting retrieved content, or misusing internal knowledge. Therefore, ensuring the trustworthiness of RAR outputs has become a critical challenge.
Uncertainty quantification (UQ) is a widely studied task in machine learning, aimed at assessing the reliability of model outputs by measuring the degree of uncertainty (or lack of confidence)  [^1] The terms uncertainty and confidence are used interchangeably in existing literature. a model has in its predictions (Malinin and Gales, 2021; Kadavath and others, 2022; Kuhn et al., 2023; Hou et al., 2024; Ye et al., 2024; Harrison et al., 2024). Recent methods of estimating the uncertainty of LLM outputs are designed for settings where the input consists solely of a query, meaning the LLM itself is the only source of uncertainty (Bakman et al., 2024; Yaldiz et al., 2025a; Duan et al., 2024). The limited work on UQ for RAG (Soudani et al., 2025; Perez-Beltrachini and Lapata, 2025) incorporates the document–response relationship into the uncertainty score, but these methods are only applicable for simple RAG settings, where documents are retrieved once and inserted into the input prompt for generation. As a result, existing UQ approaches are suboptimal for RAR.
A fundamental reason for the relatively poor performance of existing UQ methods for RAR models is that they primarily attribute uncertainty to the LLM’s generative process; i.e., next-token prediction. In RAR systems, however, we have more sources of uncertainty: retriever, which may provide irrelevant or partially relevant retrieved documents and potentially mislead the model’s reasoning and response generation processes; and generation, where the model’s reasoning may deviate from the user query’s intent and retrieved documents, leading it to formulate new search queries that fail to gather informative evidence.
In this paper, we propose Retrieval-Augmented Reasoning Consistency ($\text{R}^{2}$C), a novel UQ method that, unlike previous approaches, accounts for multiple sources of uncertainty in RAR. The central idea of $\text{R}^{2}$C is to allow the model to explore diverse reasoning paths, queries, and documents and then measure the consistency of the resulting final answers. This is achieved by modeling RAR as a Markov Decision Process (MDP) and perturbing this process in a controlled way through a set of perturbation actions across various states. Three perturbation actions are designed to influence query generation, document retrieval, and LLMs’ thinking process. These perturbations enable models to arrive at diverse final responses for uncertain generations. The confidence score is then obtained by measuring consistency of the generated answers using majority voting; see Figure 1.
We conduct our experiments across multiple datasets and five RAR models. Our experiments show that $\text{R}^{2}$C significantly outperforms existing LLM-specific UQ methods, achieving on average more than a 5% improvement in AUROC compared to the state-of-the-art UQ methods. We further extrinsically evaluate $\text{R}^{2}$C on two downstream tasks: (i) Abstention: [^2] Also referred to as selective prediction in the literature (Duan et al., 2025; Xin et al., 2021). the task of generating ‘I don’t know’ when the model is uncertain about its output (Feng et al., 2024; Madhusudhan et al., 2025), and (ii) Model Selection: [^3] Also referred to as selection-based model ensemble in the literature (Jiang et al., 2023a). the task of selecting a final answer from a pool of candidates generated by multiple systems (Chen et al., 2025c; Guha et al., 2024). Our experimental results indicate that $\text{R}^{2}$C delivers statistically significant gains over existing approaches: in Abstention, it achieves roughly 5% improvements on both F1Abstain and AccAbstain; in Model Selection, it increases exact match by about 7% relative to single RAR models and by about 3% compared to selection model baselines.
Given the strong performance of $\text{R}^{2}$C in both direct evaluation and extrinsic evaluation on downstream tasks, we investigate the factors that contributes to its effectiveness. We show that $\text{R}^{2}$C retrieves on average 25 unique documents for each score, compared to 16 documents retrieved by other UQ methods. It also achieves a query diversity of $0.35$ compared to $0.30$ other methods, measured by the inverse of the average pairwise cosine similarity between queries (Cox et al., 2021; Zhang et al., 2025). This diversity in queries and documents demonstrates that $\text{R}^{2}$C generates diverse, yet relevant reasoning paths through our controlled perturbation mechanism. As a result, this enables the method to achieve confidence scores comparable to baseline approaches while requiring only about 3 generations on average, 2.5 times fewer token generations than the 10 used by the baselines. This highlights that $\text{R}^{2}$C is not only the most effective method of its kind but also a relatively more efficient UQ approach. To summarize, the main contributions of this paper are:
  1. (1)
    We propose $\text{R}^{2}$C, a novel theocratically grounded UQ method based on MDP; the first of its kind that captures different sources of uncertainty in RAR.
  2. (2)
    We conduct extensive experiments on three dataset and five RAR methods, demonstrating the superiority of the proposed method on the UQ task with average AUROC of 82%.
  3. (3)
    We show the effectiveness of our method on both model selection and abstention tasks, significantly outperforming baselines by at least 3%.
  4. (4)
    We demonstrate that $\text{R}^{2}$C achieves an improvement in token efficiency by approximately 2.5 times.
  5. (5)
    We show that diverse query and document generation strengthens UQ by capturing multiple uncertainty sources.
2. Related Work
Retrieval-Augmented (Reasoning) Models. RAG is a framework that combines the strengths of retrieval models and generative models (Cuconasu et al., 2024). Broadly, RAG can be implemented in different ways. In the retrieve-then-generate paradigm, relevant documents are first retrieved based on the user’s input and then incorporated into the model’s prompt (Mallen et al., 2023; Soudani et al., 2024a). In contrast, Active RAG allows retrieval to occur throughout the generation process, either in fixed intervals or dynamically, whenever additional information is needed (Trivedi et al., 2023; Jiang et al., 2023b; Su et al., 2024). Retrieval-Augmented Reasoning (RAR) is a recent extension of RAG that integrates retrieval with reasoning, aiming to improve the interaction between LLMs and retrievers (Sun et al., 2025; Tran et al., 2025). For example, SelfAsk (Press et al., 2023) decomposes complex questions into follow-up queries and intermediate answers, while ReAct (Yao et al., 2023) defines a set of actions, such as search, look up, and finish, to structure interactions with external resources. More recent and effective models such as ReSearch (Chen et al., 2025a) and Search-R1 (Jin et al., 2025) are explicitly trained to seamlessly integrate external resources into reasoning. Despite their effectiveness on complex queries, RAR models remain prone to errors, such as retrieving irrelevant documents, misinterpreting content, or misusing internal knowledge.
Uncertainty Quantification (UQ) and Confidence Estimation (CE). UQ and CE are closely related but conceptually distinct. In traditional machine learning, uncertainty is defined as a property of the model’s predictive distribution given a specific input (Lin et al., 2024; Liu et al., 2025). Confidence, by contrast, reflects the model’s belief in the correctness of a particular prediction for the given input. In other words, CE is defined as the task of quantifying how certain a model is about a specific generated response (Liu et al., 2025), while UQ is a task of capturing the degree of variability or unpredictability in the model’s outputs (irrespective of a specific response) (Kendall and Gal, 2017; Gawlikowski et al., 2023). Despite this distinction, existing work sometimes use uncertainty to refer to estimated confidence (Kadavath and others, 2022; Liu et al., 2025). In this paper, we compute confidence scores for LLM generated responses. Nevertheless, following common practice in the literature (Kadavath and others, 2022; Malinin and Gales, 2021; Wang et al., 2024b; Ren et al., 2022; Fadeeva et al., 2024; Yin et al., 2024; Zhang and Zhang, 2025; Becker and Soatto, 2024; Liu et al., 2024), we refer to this confidence score as an uncertainty measure and compare it against other UQ methods.
Broadly, existing UQ methods can be divided into two categories: white-box approaches, which leverage token-level probabilities and entropy (Kadavath and others, 2022; Kuhn et al., 2023; Bakman et al., 2024; Yaldiz et al., 2025a; Duan et al., 2024; Soudani, 2025), and black-box approaches, which rely only on the final textual outputs (Lin et al., 2024; Tian et al., 2023). Most UQ methods focus on question answering and view the LLM as the only source of uncertainty, but in RAG, the retriever also contributes its own uncertainty, making this assumption incomplete. Limited research has explored UQ for RAG by modeling the document-response link, either via axioms (Soudani et al., 2025) or utility models (Perez-Beltrachini and Lapata, 2025). Some path-based approaches focus on assessing the consistency of reasoning paths in reasoning tasks (Wang et al., 2023; Qi et al., 2025; Li et al., 2025a). However, these methods do not extend naturally to RAR that involves repeated retrieval during reasoning. Recent work has studied uncertainty propagation in multi-step decision-making by combining uncertainties from intermediate steps. SAUP (Zhao et al., 2025) learns aggregation weights to merge per-step uncertainties, but it relies on ground-truth labels from the test domain. In contrast, we propose $\text{R}^{2}$C, a method that accumulates uncertainty over the entire reasoning path, while considering different sources of uncertainty including the retriever and the generator.
3. Preliminaries
Algorithm 1 $\text{R}^{2}$C: Retrieval-Augmented Reasoning Consistency
1:user query $x$, backbone LLM $\pi_{\theta}$, number of generations $B$, set of main actions $A=\{a_{ret},a_{ans}\}$, set of perturbation actions $A^{*}=\{a_{qp},a_{cr},a_{av}\}$
2:Confidence score $C(x,r)$
3:for $t=1$ to NN do $s_{t}\leftarrow\pi_{\theta}\!\left(s_{t-1},a_{t-1}\right)$ $\triangleright$ Generate the most-likely path iteratively
[…]
6:for $b=1\to B$ do
7:  $a^{*}\sim\mathcal{U}(\mathcal{A}^{*})$ $\triangleright$ Sample an action from $\mathcal{A}^{*}$
8:  if $a^{*}=a_{av}$ then $s_{t}\leftarrow s_{N}$ $\triangleright$ Select the last state for $a_{av}$
9:  else $s_{t}\sim\mathcal{U}(s_{1},s_{N-1})$ $\triangleright$ Sample a state
10:  end if
11:  $s_{t+1}=\pi_{\theta}(s_{t},a^{*})$ $\triangleright$ Apply action $a^{*}$ at state $s_{t}$
12:  for $i=1$ to $N^{b}$ do $s_{t+i+1}\leftarrow\pi_{\theta}\!\left(s_{t+i},a_{t+i}\right)$
13:  $r^{b}\leftarrow s_{N^{b}}$ $\triangleright$ Capture the sampled response
14:  $R_{G}=R_{G}\cup\{r^{b}\}$ $\triangleright$ Update the sampled response set
15:end for
[…]
17:return $c$
RAR as Markov Decision Process. We formalize RAR as a stochastic Markov Decision Process (MDP), described by a quadruple $(S,A,P,R)$, where $S$ denotes a set of states, $A$ represents a set of actions the agent can take in a state, $P(s_{t+1}\mid s_{t},a_{t})$ denotes the probability of transitioning from state $s_{t}$ to state $s_{t+1}$ given action $a_{t}$, and $R(s_{t},a_{t})$ is the reward received by the agent after taking action $a_{t}$ in state $s_{t}$. To generate a factual response to a user query $x$, the agent $\pi$ starts from an initial state $s_{0}$ corresponding to the query $x$. The agent then iteratively selects an action at each step until it chooses a halting action and generates the response. The agent assigns a probability $p_{\pi}(a_{t}\mid s_{t})$ to each possible action based on the current state $s_{t}$.
In our unified formulation of MDP for RAR, the LLM $\pi_{\theta}$ acts as the agent. The environment can take various forms, such as a knowledge repository (Yao et al., 2023). The set of possible actions in each state is $A=\{a_{\text{ret}},a_{\text{ans}}\}$, where $a_{\text{ans}}$ denotes the halting action, and $a_{\text{ret}}$ represents the retrieval action. Each intermediate state $s_{t}$ consists of a think $\tau_{t}$ followed by a search query $q_{t}$; i.e., $s_{t}=\langle\tau_{t},q_{t}\rangle$. The final state $s_{N}$ contains a think $\tau_{N}$ followed by a final response $r$. The transition probability $P(s_{t+1}\mid s_{t},a_{t})$ is determined by the LLM itself. An explicit reward function is not always required, as some agents operate without additional training.
Consistency-based Uncertainty Quantification. The core idea of consistency-based methods (Xiong et al., 2023; Hou et al., 2024) is to generate multiple responses for a given query by varying either the input prompt or the temperature parameter used in stochastic decoding (Korikov et al., 2025). The idea build upon the hypothesis of self-consistency theory, where correct reasoning processes, even if they are diverse, tend to have greater agreement in their final answer (Wang et al., 2023). The pairwise similarity among these responses is then computed and aggregated into a single confidence score (Lin et al., 2024; Bakman et al., 2025). Formally, consider a model $\pi$ parameterized with $\theta$, generates the most-likely response $r$ by setting the sampling temperature to less than one. Then, $B$ additional responses $R_{G}=\{r^{b}\}_{b=1}^{B}$ are sampled using various sampling strategies such as increasing the temperature (Wang et al., 2023), changing the input (Korikov et al., 2025), or altering the reasoning path (Qi et al., 2025; Li et al., 2025a). A transformation function $\mathcal{\phi}$ is then applied to convert $R_{G}$ into a confidence score. Our method builds on the reasoning path perturbation approach and employs Majority Voting (Wang et al., 2023) as the transformation function, where confidence is the degree of consistency, measured by the proportion of sampled responses that match the most-likely response:
(1) $C(x,r)=\mathcal{\phi}(R_{G},r)=\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}(r^{b}\equiv r).$
4. Methodology
We propose Retrieval-Augmented Reasoning Consistency, $\text{R}^{2}$C, to address UQ in RAR models. $\text{R}^{2}$C is a consistency-based approach that performs UQ in two main stages, as illustrated in Figure 1: (i) generating the most-likely response, and (ii) sampling multiple generations. The core idea of $\text{R}^{2}$C is to perturb the reasoning paths of these multiple generations through a set of perturbation actions, denoted as $A^{*}$. In MDP terms, $\text{R}^{2}$C temporarily replaces the main action set $A$ with $A^{*}$ for a single state, allowing the RAR model to interleave its generation process and explore new reasoning trajectories, queries, and documents. We define three perturbation actions employed in $\text{R}^{2}$C: (i) Query Paraphrasing, $a_{qp}$, (ii) Critical Rethinking, $a_{cr}$, and (iii) Answer Validation, $a_{av}$. By employing a variety of actions to perturb the reasoning processes of LLMs, $\text{R}^{2}$C captures both epistemic (model) uncertainty and aleatoric (data) uncertainty. This perspective is supported by Liu et al. (2025), who argue that reasoning uncertainty encompasses both epistemic and aleatoric components. In this work, we do not seek to disentangle these two sources of uncertainty; rather, we measure their combined effect as total uncertainty.
We note at the outset that $\text{R}^{2}$C builds on the following definition for perturbation actions:
Definition. A perturbation action changes the state of a system without omitting existing information or adding extra knowledge to the system.
This aligns our method with the self-consistency theory and ensures that any divergence between the most-likely response and the generated response after applying a perturbation action is attributable to the variability of stochastic response generation, rather than to changes in the system’s knowledge. In the following sections, we first formally describe how $\text{R}^{2}$C perturbs the generation path and then describe perturbation actions.
4.1. $\text{R}^{2}$C: Retrieval-Augmented Reasoning Consistency
$\text{R}^{2}$C models uncertainty for RAR as an MDP, in which multiple response generations are produced by temporarily replacing the action set $A$ with an alternative set $A^{*}$ at a randomly selected state $s_{t}$. First, the most-likely generation is produced as an MDP iteratively:
$s_{t}\leftarrow\pi_{\theta}\!\left(s_{t-1},a_{t-1}\right);\quad t=1,\dots,N,$
where $N$ denotes the length of the reasoning path, determined by the agent $\pi_{\theta}$ when it selects the halting action $a_{\text{ans}}$. The most-likely response $r$ is obtained from the final state $s_{N}$. For example, in the left part of Figure 1, the most-likely response is “Orlando Magic,” derived from the most-likely generation highlighted in yellow. The middle part of the figure shows the multi-generation process where an action is randomly selected and applied to a state.
To construct the sampled response set $R_{G}$, we first fix the number of generations to $B$. In each generation, we uniformly sample an action $a^{*}$ from the perturbation action set $A^{*}$ and a perturbation state $s_{t}$ from the most-likely reasoning path $\{s_{t}\}_{t=1}^{N}$. An exception is the action $a_{av}$, for which the perturbation state is always set to $s_{N}$. The agent then transitions from state $s_{t}$ to $s_{t+1}$ given action $a^{*}$:
$s_{t+1}\leftarrow\pi_{\theta}(s_{t},a^{*}).$
For the remainder of the path, the actions are sampled from the main action set $A$ until reaching a new end state $N^{b}$, determined by the agent:
$s_{t+i+1}\leftarrow\pi_{\theta}\!\left(s_{t+i},a_{t+i}\right);\quad i=1,\dots,N^{b}.$
The final sampled response $r^{b}$ is obtained from the last state $s_{N^{b}}$ and added to $R_{G}$. After $B$ iterations, we obtain the response set $R_{G}=\{r^{b}\}_{b=1}^{B}$, on which the confidence score function (Eq. (1)) is applied to compute the final confidence score. Algorithm 1 provides a detailed description of the entire process in $\text{R}^{2}$C.
4.2. Perturbation Actions
A1: Query Paraphrasing (QP). Constructing effective search queries is critical for retrieving relevant documents; however, LLMs are not inherently optimized for this purpose (Jiang et al., 2025b; Ma et al., 2023; Li et al., 2025a). The QP action ($a_{qp}$) is introduced as a query optimization mechanism that enables the system to explore alternative semantic formulations of the original query. Precisely, when action $a_{qp}$ is applied, the think $\tau_{t}$ of the state $s_{t}$ is preserved and only the query changes. Formally, the LLM $\pi_{\theta}$ takes action $a_{qp}$ at state $s_{t}$, transitioning to state $s_{t+1}$ with the same think $\tau_{t}$ and a new query $q_{t+1}$:
$s_{t+1}=\langle\tau_{t},q_{t+1}\rangle\leftarrow\pi_{\theta}(a_{qp},\langle\tau_{t},q_{t}\rangle).$
Conceptually, QP perturbation tests whether the reasoning path is so fragile that paraphrasing the search query can alter its direction and lead to the retrieval of different documents. The QP action is implemented by prompting the LLM with a paraphrasing instruction.
Table 1. Performance of UQ methods measured by AUROC. In each column, the best and second-best methods are indicated by bold and underline, respectively. Superscripts and denote statistically significant differences according to the DeLong test ($p$ ¡ 0.05), compared to ReaC and P(true), respectively, which are the two best-performing methods on average.
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) ReSearch (Chen et al., 2025a) Search-R1 (Jin et al., 2025) Avg.
Uncer. M. Popqa Hotp. Musi. Popqa Hotp. Musi. Popqa Hotp. Musi. Popqa Hotp. Musi. Popqa Hotp. Musi.
PE (Kadavath and others, 2022) 55.34 59.11 48.61 36.75 39.95 41.14 39.86 53.44 51.59 40.93 61.03 56.69 48.49 48.27 50.90 48.81
SE (Kuhn et al., 2023) 64.26 68.01 54.68 49.73 40.37 41.69 67.56 65.84 54.61 53.88 64.72 63.43 59.11 53.45 61.38 57.51
MARS (Bakman et al., 2024) 54.70 59.48 51.53 41.29 40.42 41.97 43.28 57.69 56.80 40.03 61.76 57.58 48.33 49.28 49.94 50.27
SAR (Duan et al., 2024) 51.65 63.05 52.31 29.56 32.90 31.54 40.92 52.37 44.66 41.67 62.75 51.47 45.87 46.40 45.22 46.16
LARS (Yaldiz et al., 2025a) 73.97 70.03 66.45 79.95 68.28 71.87 83.79 71.42 66.19 76.95 66.12 71.64 71.54 67.08 71.62 71.79
NumSS (Kuhn et al., 2023) 71.31 65.19 62.75 74.90 63.73 62.59 78.88 63.74 62.12 80.42 64.46 69.34 69.76 64.10 65.72 67.93
EigV (Lin et al., 2024) 70.53 66.80 62.13 76.48 66.44 57.44 79.72 68.58 64.75 80.63 66.43 66.31 69.33 66.54 64.27 68.43
ECC (Lin et al., 2024) 72.89 69.95 64.11 80.27 69.51 61.92 81.98 69.27 69.65 81.85 68.49 72.55 70.87 67.76 67.65 71.25
Deg (Lin et al., 2024) 70.53 66.79 61.77 76.70 67.70 58.12 80.69 68.87 66.29 81.67 67.15 67.85 69.51 66.75 64.53 68.99
RrrC (Li et al., 2025a) 71.14 71.17 81.28 48.02 68.30 75.99 65.95 73.87 77.95 68.30 71.63 74.25 68.92 71.08 70.77 70.57
SelfC (Wang et al., 2023) 74.33 69.06 68.40 80.73 75.34 72.96 81.26 72.34 76.87 82.01 72.14 77.89 71.63 69.04 70.36 74.29
ReaC (Qi et al., 2025) 77.02 70.90 76.75 81.53 76.94 74.74 81.09 74.97 77.01 82.50 75.22 77.86 73.22 72.79 75.29 76.52
P(true) (Kadavath and others, 2022) 76.51 77.66 78.65 75.07 73.45 71.51 78.57 73.19 74.83 84.35 77.77 81.42 75.73 76.25 74.80 76.65
$\text{R}^{2}$C(our) 80.08 81.09 75.82 84.75 83.25 81.16 87.09 79.66 83.22 86.02 80.76 82.39 84.92 79.51 80.08 81.99
A2: Critical Rethinking (CR). RAR models often suffer from the problem of self-criticism, where they fail to recognize that previously retrieved information is noisy or irrelevant (Jin et al., 2025; Jiang et al., 2025b; Asai et al., 2024; Li et al., 2025a). Consequently, they continue to build their reasoning on top of earlier steps, even when those steps are uninformative and lack relevant content. This issue becomes particularly severe when it occurs in the early stages of the reasoning path.
The CR action ($a_{cr}$) critically reassesses the reasoning states produced up to state $s_{t}$. When applied at state $s_{t}$, it introduces a new state $s_{t+1}$ that the think $\tau_{t+1}$ explicitly evaluates the previously retrieved information as unhelpful, irrelevant, or misleading, and the accompanying search query $q_{t+1}$ is formulated to support this critical assessment. Formally, in state $s_{t}$, the LLM $\pi_{\theta}$ is prompted with a critical rethinking instruction to generate a new state $s_{t+1}$:
$s_{t+1}=\langle\tau_{t+1},q_{t+1}\rangle\leftarrow\pi_{\theta}(a_{cr},\langle\tau_{t},q_{t}\rangle).$
Conceptually, if the reasoning path so far has been incorrect, this action enables the system to adjust toward a more reliable trajectory. If the path has been correct, CR strengthens its validity, thereby increasing confidence in the final outcome.
A3: Answer Validation (AV). RAR models face difficulties with validating their final response, detecting whether the generated response meets certain criteria of the query (Wang et al., 2024a; Salemi et al., 2025; He et al., 2024; Jiang et al., 2025a). One challenge is that the response is built upon a reasoning path that integrates both documents and the intermediate reasoning trajectory, which often leads the LLM to exhibit excessive confidence in its output. Another challenge is that different tasks and response types involve specific validation criteria, but RAR models are generally unaware of these requirements.
We introduce the AV action ($a_{av}$) to validate the final response by prompting the LLM to reconsider its generation once a response has been produced, based on predefined criteria. Specifically, the LLM first generates a query-aware reasoning path summary (Shi et al., 2025; Li et al., 2025b; Achkar et al., 2025), and then evaluates the final response using two criteria: (i) Groundedness: is the response supported by the retrieved documents? and (ii) Correctness: does the response appropriately and sufficiently address the query, given the available evidence? Formally, let $D=\{D_{1},\dots,D_{N-1}\}$ be the set of documents retrieved at states $s_{1},s_{2},...s_{N-1}]$. A model $\mathcal{M}$ generates the summary $\hat{S}$ of these documents: $\hat{S}=\mathcal{M}(x,D).$ The state $s_{N}$ is then updated with this summary, denoted as $\hat{s}_{N}$. With the updated state $\hat{s}_{N}$, the LLM $\pi_{\theta}$ is then instructed to generate a new state $s_{N+1}$:
$s_{N+1}\leftarrow\pi_{\theta}(a_{av},\hat{s}_{N}).$
In principle, if the final response $r$ is validated as correct, the system outputs it directly in state $s_{N+1}$. Otherwise, if the validation indicates that the answer is incorrect or incomplete, the system begins a new reasoning path starting from $\langle\tau_{N+1},q_{N+1}\rangle$.
5. Experimental Setup
Our experiments consist of evaluation of uncertainty scores estimated by UQ methods as well as extrinsic evaluation on Abstention and Model Selection. In the following, we review our experimental setup for each of these tasks.
5.1. Direct Evaluation of UQ Estimations
Datasets: We evaluate $\text{R}^{2}$C on both single-hop and multi-hop QA tasks using the PopQA (Mallen et al., 2023), HotpotQA (Yang et al., 2018), and Musique (Trivedi et al., 2022) datasets. Following prior work (Trivedi et al., 2023; Jiang et al., 2023b; Yao et al., 2023; Moskvoretskii et al., 2025; Bakman et al., 2025), we randomly sampled $500$ queries from each dataset as the test set. We will release our sampled queries for improving reproducibility of our work. For the retrieval corpus, we use the 2018 Wikipedia dump (Karpukhin et al., 2020). [^4] https://huggingface.co/datasets/PeterJinGo/wiki-18-corpus The number of retrieved documents is fixed to three across all models (Mallen et al., 2023; Jin et al., 2025; Yao et al., 2023; Press et al., 2023).
Evaluation Metrics: To evaluate the quality of outputs, we follow Jin et al. (2025) and report the exact match, where a prediction is counted as correct if and only if it exactly matches one of the ground-truth responses. For evaluating UQ methods, we follow prior work on UQ and use the threshold-free metric AUROC, which captures the correlation between uncertainty scores and response correctness (Kadavath and others, 2022; Kuhn et al., 2023; Bakman et al., 2024). As suggested by Perez-Beltrachini and Lapata (2025), significant differences between two AUROC values are assessed using the paired De Long test (DeLong et al., 1988).
Table 2. Abstention performance measured by AbstainAccuracy and AbstainF1 at a 0.9 confidence threshold. For each column, the best and second-best methods are indicated in bold and underlined, respectively. A superscript denotes a statistically significant difference compared to ReaC based on the McNemar test for Accuracy and the Bootstrap test for F1 ($p<0.05$).
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) ReSearch (Chen et al., 2025a) Search-R1 (Jin et al., 2025) Avg.
Uncer. M. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq.
Abstain Accuracy
RrrC (Li et al., 2025a) 61.4 65.6 73.0 60.8 72.2 82.4 56.6 66.2 68.2 59.2 64.4 64.0 60.6 64.0 63.6 65.48
P(true) (Kadavath and others, 2022) 70.6 70.8 80.4 70.4 67.2 75.0 72.6 65.8 75.0 78.8 71.4 78.2 70.8 70.0 74.4 72.76
SelfC (Wang et al., 2023) 68.6 64.2 76.2 74.0 75.8 87.6 75.0 69.0 90.2 75.2 68.6 83.6 65.4 62.8 77.8 74.27
ReaC (Qi et al., 2025) 69.2 64.6 80.4 74.4 75.4 89.4 75.6 71.4 87.4 77.2 68.6 80.8 67.8 68.8 80.6 75.44
$\text{R}^{2}$C (our) 77.2 74.4 88.6 77.0 76.8 90.4 80.4 74.4 89.4 81.6 74.8 84.0 77.0 73.4 84.4 80.25
Abstain F1
RrrC (Li et al., 2025a) 62.52 70.54 82.75 71.59 80.82 89.69 56.33 71.50 78.99 54.86 63.37 73.68 55.12 61.20 73.92 69.79
P(true) (Kadavath and others, 2022) 73.60 75.33 88.27 72.07 73.46 84.58 76.50 71.73 84.51 81.20 73.46 85.82 71.03 71.15 83.37 77.74
SelfC (Wang et al., 2023) 72.60 69.51 85.71 79.10 83.80 93.18 80.50 78.07 94.64 81.71 75.19 90.57 62.31 62.34 86.51 79.72
ReaC (Qi et al., 2025) 73.54 70.45 88.41 79.01 83.17 94.19 81.24 79.48 93.03 81.25 72.60 88.43 67.60 71.11 88.27 80.79
$\text{R}^{2}$C (our) 82.35 81.76 93.77 83.69 85.08 94.81 86.42 82.22 94.27 84.40 78.71 90.49 80.80 77.57 90.97 85.82
Models: In line with prior work in RAR (Jin et al., 2025; Chen et al., 2025a), we employ Qwen-2.5-7B-Instruct (Yang and others, 2024) as the generator LLM and path summary generator for action $a_{av}$. For UQ, we sample $10$ responses per query with a temperature of $T=1.0$, while for correctness evaluation we generate the most-likely generation with $T=0.7$ (Soudani et al., 2025; Bakman et al., 2024). Retrieval is performed using a two-stage re-ranking pipeline: BM25 (Robertson and Walker, 1994) is used for initial retrieval, followed by re-ranking with the pre-trained cross-encoder model ms-marco-MiniLM-L-6-v2 from the sentence-transformers library. [^5] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2 All experiments are conducted on four Nvidia A100 GPU, each with 40 GB memory, requiring $\sim$1500 GPU hours in total.
Baselines: We use two sets of baselines: (1) Path-based methods, which focus on generating multiple responses based on diverse reasoning paths and differ mainly in how they initiate new generations relative to the most-likely generation. Self-Consistency (SelfC) (Wang et al., 2023) ignores the most-likely generation and instead produces diverse set of reasoning paths, starting from scratch. Reasoning Consistency (ReaC) (Qi et al., 2025) randomly truncates the most-likely reasoning path at different random steps and regenerates the response based on the subsequent reasoning steps. Retrieval-Retained Reasoning Consistency (RrrC) (Li et al., 2025a) applies truncation only after the last retrieved document. In all cases, the generated responses are aggregated into a final score using majority voting. 2) Estimation-based methods include both white-box and black-box approaches. The white-box methods are PE (Kadavath and others, 2022), SE (Kuhn et al., 2023), MARS (Bakman et al., 2024), LARS (Yaldiz et al., 2025a), and SAR (Duan et al., 2024). The black-box methods are NumSS (Kuhn et al., 2023), EigV, ECC, Deg (Lin et al., 2024), and P(true) (Kadavath and others, 2022). All of these methods rely on generations obtained in the same way as SelfC. All UQ methods are implemented using TruthTorchLM (Yaldiz et al., 2025b).
5.2. Extrinsic UQ Evaluation via Abstention
Task Formulation. Declining to respond due to uncertainty is an important application of UQ (Feng et al., 2024). An abstention function determines whether the model should withhold an answer. In our setup, this decision is guided by confidence scores. We introduce a threshold $\tau_{\text{abs}}$: if the confidence is less than $\tau_{\text{abs}}$, the model abstains; otherwise, it produces an answer. Formally, given a response $r$ to a query $x$ and a confidence function $C$, we define the abstention function $f_{\text{abs}}$ as:
$f_{\text{abs}}(r)=\begin{cases}\text{true},&\text{if }C(x,r)<\tau_{\text{abs}}\\ \text{false},&\text{otherwise}.\end{cases}$
Baselines & Evaluation Metrics. As baselines, we implement the abstention task using confidence scores derived from different UQ methods, including P(true), RrrC, ReaC, SelfC, and $\text{R}^{2}$C. For evaluation, we follow Feng et al. (2024) and report two metrics: AbstainAccuracy and AbstainF1. These metrics assess how well a model balances answer correctness with appropriate abstention. We define a confusion matrix with four outcomes: (A) answered correctly, (B) abstained correctly, (C) answered incorrectly, and (D) abstained incorrectly. Based on this matrix, AbstainAccuracy is defined as $\frac{A+D}{A+B+C+D}$ and measures whether abstention decisions are correct overall. Ideally, a model should abstain when it would otherwise answer incorrectly, and provide an answer when it is correct. AbstainF1 captures the trade-off between reliability and answer coverage. It is computed as the harmonic mean of precision ($\frac{D}{B+D}$) and recall ($\frac{D}{C+D}$). This metric penalizes both unnecessary abstentions and incorrect answers, providing a balanced evaluation of abstention behavior. The threshold $\tau_{\text{abs}}$ for each UQ method is selected by performing sweep using validation sets.
5.3. Extrinsic UQ Evaluation via Model Selection
Task Formulation. Model selection (or selection-based model ensemble (Chen et al., 2025c; Jiang et al., 2023a)) aims to select a final response for a question based on multiple candidate responses generated by different systems (Chen et al., 2025c). Formally, given a user query $x$ and a set of systems $\{S_{1},S_{2},\dots,S_{M}\}$, each system $S_{i}$ produces a response candidate $r_{i}$. A model selection method $\mathcal{M}(x,R)$ then selects the final response $\hat{r}$ from all candidate responses $R=\{r_{1},r_{2},\dots,r_{M}\}$.
$\text{R}^{2}$C Select. The proposed $\text{R}^{2}$C Select utilizes the confidence scores derived from $\text{R}^{2}$C. Our method groups semantically similar responses into $K$ clusters. Following (Li et al., 2025a; Yaldiz et al., 2025b), we use Qwen-2.5-7B-Instruct to compute pairwise semantic similarities between candidate responses and cluster similar ones together. We then assign a confidence score to each cluster $g_{i}$:
$R_{G}=\{\langle r_{{}_{1}},c_{g_{1}}\rangle,\langle r_{{}_{2}},c_{g_{1}}\rangle,\dots,\langle r_{{}_{K}},c_{g_{K}}\rangle\}.$
The confidence score of each cluster $g=\{r_{i}\}_{i=1}^{m}$ is computed by aggregating the confidence scores of its members: $c_{g}=\sum_{r_{i}\in g}C(x,r_{i}),$ where the $C$ function provides the confidence score for the response $r_{i}$. The final response $\hat{r}$ is the response with the highest confidence score: $\hat{r}=\arg\max_{j}c_{g_{j}}.$ If no clustering is applied, each response constitutes a cluster and $M=K$; we refer to this variation as $\text{R}^{2}$C Select w/o clustering.
Table 3. Model Selection performance measured by exact match. The superscript denotes a statistically significant difference from the best-performing baseline (underlined), according to the Wilcoxon test ($p$ ¡ 0.05).
RAG System PopQA HotpotQA Musique Average
Vanilla LLM & RAG
Direct 18.8 20.8 2.6 14.1
CoT 17.6 22.2 5.8 15.2
Vanilla RAG 30.2 18.6 4.4 17.7
IRCoT (Trivedi et al., 2023) 34.6 27.2 5.4 22.4
FLARE (Jiang et al., 2023b) 31.6 25.2 8.4 21.7
DRAGIN (Su et al., 2024) 28.6 23.2 4.6 18.8
Retrieval-Augmented Reasoning (RAR)
SelfAsk (Press et al., 2023) 35.6 33.0 10.4 26.3
ReAct (Yao et al., 2023) 36.8 27.8 10.8 25.1
Search-o1 (Li et al., 2025b) 33.2 29.0 10.0 24.1
ReSearch (Chen et al., 2025a) 38.6 38.8 16.6 31.3
Search-R1 (Jin et al., 2025) 41.6 41.4 16.0 33.0
Model Selection RAR
Random 31.4 31.6 10.6 24.5
LLMBlender (Jiang et al., 2023a) 34.4 36.0 12.0 27.5
- w/o clustering 35.6 31.8 9.6 25.7
- w/o clus. & conf. 32.0 26.0 8.6 22.2
RAGEnsemble (Chen et al., 2025b) 45.6 46.0 19.4 37.0
- w/o clustering 44.4 40.8 18.0 34.4
- w/o clus. & conf. 43.2 37.2 13.0 31.1
$\text{R}^{2}$CSelect (our) 46.8 50.2 22.6 39.9
- w/o clustering 45.4 44.0 19.2 36.2
Ideal Model Selection 55.0 57.2 30.0 47.4
Baselines & Evaluation Metrics. We evaluate $\text{R}^{2}$C Select performance against both single RAG and RAR systems, as well as existing selection-based ensemble approaches. As baselines, we consider LLMBlender (Jiang et al., 2023a) and RAGEnsemble (Chen et al., 2025b). LLMBlender is a trained reward model that given a user query and a set of candidate responses, ranks the responses accordingly. RAGEnsemble is an instruction-based approach that selects a single final answer from a set of candidate responses. For all selection methods, including the baselines and $\text{R}^{2}$C Select, the response candidates are obtained from SelfAsk (Press et al., 2023), ReAct (Yao et al., 2023), Search-o1 (Li et al., 2025b), ReSearch (Chen et al., 2025a), and Search-R1 (Jin et al., 2025). For evaluation, we report the correctness of the final answer using exact match.
6. Results
We present a set experiments that address the following research questions: RQ1: How does $\text{R}^{2}$C perform in quantifying uncertainty for different RAR models? (Sec. 6.1), RQ2: How does $\text{R}^{2}$C perform as an external signal on downstream tasks, such as Abstention and Model Selection? (Sec. 6.2), RQ3: What factors contribute to the effectiveness of $\text{R}^{2}$C? (Sec. 6.3), RQ4: How does $\text{R}^{2}$C balance effectiveness and efficiency? (Sec. 6.4), RQ5: What is the effect of different actions in the performance of $\text{R}^{2}$C? (Sec. 6.5). RQ6: Does $\text{R}^{2}$C generalize across different backbone LLMs? (Sec. 6.6)
6.1. Uncertainty Quantification Performance
RQ1 evaluates the performance of $\text{R}^{2}$C compared to other UQ methods. Table 1 presents results on five RAR systems across three datasets. The findings indicate that most white-box methods, i.e., PE (Kadavath and others, 2022), SE (Kuhn et al., 2023), MARS (Bakman et al., 2024), and SAR (Duan et al., 2024), perform relatively poorly, with AUROC values ranging from about 30 to 60. This weakness stems from their overreliance on token probabilities. In contrast, black-box methods, such as NumSS (Kuhn et al., 2023), EigV, ECC, and Deg (Lin et al., 2024), generally outperform white-box methods, reaching average AUROC values between roughly 60 and 80. Interestingly, P(true) ranks as the second-best method in terms of average AUROC across all approaches, highlighting that black-box methods become much more effective when using an LLM as the scoring function. This advantage largely comes from their stronger reliance on textual diversity. The supervised method LARS, optimized for QA with no reasoning, achieves AUROC scores between 65 and 85 across all cases, surpassing both white-box and black-box approaches on average. This finding highlights the potential of as a promising direction for future research.
Figure 2. Distribution of the number of unique retrieved documents for the most-likely path (main) and multi-generations.
Figure 3. Distribution of diversity scores for search queries generated for each user query across reasoning paths.
Path-based methods are another group of approaches that we evaluate, including the proposed $\text{R}^{2}$C. RrrC achieves an average AUROC of 70.57, which is the lowest among these methods. This indicates that simply keeping the documents in the reasoning path and regenerating the last state is not effective. SelfC and ReaC perform relatively on par, and ReaC outperforms SelfC on average, showing that regenerating from the top of the reasoning path does not necessarily guarantee a better uncertainty score. Finally, $\text{R}^{2}$C outperforms all methods by a large margin, with an absolute improvement of 5% on average. These findings suggest that applying actions to the reasoning path enables exploration of a wider variety of possible reasoning states, leading to more reliable scores.
Figure 4. Performance of UQ methods with varying numbers of generations.
Figure 5. Performance of UQ methods with varying number of generated tokens, illustrating the trade-off between effectiveness and efficiency.
6.2. Extrinsic Evaluation Results
RQ2 evaluates the performance of $\text{R}^{2}$C in Abstention and Model Selection. For Abstention, Table 2 shows the performance on RAR models and datasets. Abstain Accuracy measures whether both abstentions and non-abstentions are detected correctly, while captures the effectiveness of detecting abstain samples. Considering Abstain Accuracy, $\text{R}^{2}$C outperforms the baselines in all cases except on the Musique dataset and performs on par with the second-best baseline in ReAct, Search-o1, and ReSearch.Considering Abstain F1, $\text{R}^{2}$C significantly outperforms the best baseline in all setup, expect for ReAct on Musique. This indicates that $\text{R}^{2}$C scores are more reliable in detecting than correct abstentions than correct non-abstentions. Overall, compared to other UQ methods, $\text{R}^{2}$C achieves significantly better performance in most cases—on average, about 5% higher than the second-best model. These results demonstrate that the uncertainty scores generated by $\text{R}^{2}$C are reliable enough for the system to decide when to refrain from answering.
For model selection, Table 3 reports the results of individual RAG and RAR systems, as well as selection-based ensemble models. Among individual systems, RAR models outperform RAG models for both simpler single-hop questions (PopQA) and more complex multi-hop queries (HotpotQA and Musique). Within selection-based ensemble models, LLMBLender performs poorly, even worse than fine-tuned single RAR models such as ReSearch and Search-R1. In contrast, RAGEnsemble surpasses both LLMBLender and all single RAR models, highlighting the advantage of an instruction-tuned model for selecting the final response. Finally, $\text{R}^{2}$C Select achieves the best performance, significantly outperforming all RAR and RAG systems, with 3.7% average improvement on HotpotQA and Musique datasets. These results confirm that $\text{R}^{2}$C score is a reliable criterion, not only within a single system but also across different systems. Interestingly, using $\text{R}^{2}$C confidence scores alone, without clustering, improves selection-based ensemble models in comparable settings, highlighting informativeness of $\text{R}^{2}$C scores even for models that are not trained on confidence.
Figure 6. Performance of $\text{R}^{2}$C with different action sets.
Table 4. Performance of UQ methods measured by AUROC across different LLMs. For each LLM and for each column, the best and second-best methods are indicated by bold and underline, respectively. Superscripts and denote statistically significant differences according to the DeLong test ($p<0.05$), compared to SelfC and ReaC, respectively, which are the two best-performing methods on average.
Gemma2-9b-it GPT-4o-mini
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) Avg. SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) Avg.
Uncer. M. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq.
RrrC (Li et al., 2025a) 76.64 71.01 76.16 57.74 72.93 75.32 55.72 63.96 64.33 68.20 63.07 70.96 83.69 53.21 58.16 54.78 70.49 77.27 84.04 68.41
P(true) (Kadavath and others, 2022) 71.33 74.90 73.44 76.56 73.08 75.26 76.32 72.45 74.62 74.22
SelfC (Wang et al., 2023) 76.82 72.70 73.59 81.79 74.34 76.84 79.17 70.59 76.07 75.77 76.59 58.28 54.04 81.08 73.32 75.83 78.52 74.39 81.83 72.65
ReaC (Qi et al., 2025) 76.24 79.77 78.73 80.16 71.33 80.80 78.45 73.61 75.77 77.21 75.53 60.58 54.66 81.53 74.34 81.10 79.41 75.09 85.09 74.15
$\text{R}^{2}$C (our) 80.29 81.86 80.31 82.87 80.62 80.37 87.65 76.95 77.39 80.92 86.56 87.82 84.28 81.56 75.84 81.94 82.48 84.59 86.29 83.48
6.3. Explaining the Strength of $\text{R}^{2}$C
RQ3 examines why $\text{R}^{2}$C outperforms other UQ methods. To this end, we analyze the diversity of reasoning paths as an indicator of how effectively $\text{R}^{2}$C captures the uncertainty arising from both the generator and the retriever, i.e., the two key sources of uncertainty in our framework. We report the number of unique documents retrieved during the multi-generation step, as shown in Figure 2. On average, $\text{R}^{2}$C retrieves 24.71 unique documents for a single confidence score, whereas RrrC, SelfC, and ReaC retrieve 5.81, 15.35, and 16.4 documents, respectively.
We further report on the diversity of generated search queries, as a proxy for LLM’s thinking process. Query diversity is measured based on the average pair-wise similarity of generated queries (Cox et al., 2021; Zhang et al., 2025). Formally, given $n$ search queries $\{q_{1},\dots,q_{n}\}$, query diversity is computed as:
$\text{Query Diversity}=1-\frac{2}{n(n-1)}\sum_{i=1}^{n}\sum_{i<j}^{n}cos(e_{i},e_{j}),$
where $cos(.)$ represents cosine similarity function, and $e_{i}$ denotes the normalized embedding of query $q_{i}$ (i.e., $|e_{i}|=1$) obtained from sentence-transformers/all-MiniLM-L6-v2. Figure 3 presents the results. On average, $\text{R}^{2}$C achieves a diversity score of $0.35$, while RrrC, SelfC, and ReaC achieve $0.20$, $0.28$, and $0.30$, respectively. These results indicate that $\text{R}^{2}$C generates more diverse search queries when estimating confidence. These findings indicate $\text{R}^{2}$C effectively captures the uncertainty of both the retriever and the generator by sequentially diversifying their inputs.
6.4. Effectiveness vs. Efficiency
RQ4 explores the trade-off between effectiveness and efficiency of $\text{R}^{2}$C. Figure 4 presents the relationship between AUROC and the number of response generations across various datasets and RAR models. The results demonstrate that $\text{R}^{2}$C consistently outperforms other methods, even with a smaller number of generations. On average, across all datasets and RAR models, $\text{R}^{2}$C achieves an AUROC performance of about 77% with only three generations, which is comparable to the performance of SelfC and ReaC, requiring 10 generations to reach a similar level.
For a deeper exploration of the trade-offs between effectiveness and efficiency, we measure efficiency using the total number of generated tokens, following prior work (Perez-Beltrachini and Lapata, 2025). Figure 5 reports AUROC performance for equal numbers of generated tokens. It shows $\text{R}^{2}$C achieves higher AUROC scores than other methods given the same number of generated tokens for multi-hop datasets. For the single-hop dataset, PopQA, $\text{R}^{2}$C performs comparably to SelfAsk, ReAct, and ReSearch, but surpasses them on Search-o1 and Search-R1. Moreover, $\text{R}^{2}$C produces on average around 700 tokens with three generations, reaching the same AUROC score of baselines with approximately 1,700 tokens with 10 generations. These findings demonstrate that $\text{R}^{2}$C improves efficiency by roughly 2.5 times. Overall, these results indicate that $\text{R}^{2}$C outperforms other UQ methods in terms of both effectiveness and token generation efficiency. Beyond improving token-generation efficiency, $\text{R}^{2}$C is well suited for real-world deployment because it enables parallel generation of fully independent responses, thereby addressing the typical latency concerns associated with sampling-based methods. In practice, the confidence score can be computed at the cost of at most one additional generation beyond the most-likely response.
6.5. Action Selection
RQ5 investigates the roll of action set in $\text{R}^{2}$C. Figure 6 illustrates the performance of $\text{R}^{2}$C with different action sets. In the first three bars, only a single action is used, meaning that in each generation, the cut point in the reasoning path is chosen randomly, while the action itself remains fixed. The fourth bar represents our main setup, where both the action and the cut point are selected randomly. For Search-R1 and PopQA, the main setup outperforms all other configurations, whereas in HotpotQA, QP performs slightly better. For SelfAsk, the AV action achieves the best performance, with the main setup ranking second. These results indicate that while our main setup is generally robust, there are still potentials to design action configurations better suited to specific RAR systems. To further improve $\text{R}^{2}$C, future work could focus on learning or adaptively selecting perturbations (e.g., via reinforcement learning), which we expect would enhance both effectiveness and efficiency.
6.6. Generalization Across Backbone LLMs
RQ6 investigates the robustness of $\text{R}^{2}$C to changes in the generation model. Table 4 summarizes the performance of UQ methods under alternative LLM backbones. To ensure diversity, we evaluate one open-source LLM, Gemma2-9b-it, [^6] https://huggingface.co/google/gemma-2-9b-it and one closed-source LLM, GPT-4o-mini. We do not evaluate ReSearch and Search-R1 in this setting, as their fine-tuned models are only available for Qwen family. With Gemma2-9b-it as the backbone, $\text{R}^{2}$C outperforms the second-best method, ReaC, by more than 3%, indicating that the effectiveness of $\text{R}^{2}$C is largely independent of the choice of LLM.
For GPT-4o-mini, $\text{R}^{2}$C outperforms ReaC by more than 9%. The P(true) metric cannot be reported in this setting because token-level probabilities are not accessible. In the case of SelfAsk, we observe a different performance ordering, in which RrrC emerges as the second-best method after $\text{R}^{2}$C, most notably on the Musique dataset. This shift is primarily due to the reduced effectiveness of SelfAsk when GPT-4o-mini is used as the backbone compared to Qwen-2.5-7B-Instruct, despite additional prompt engineering. In particular, SelfAsk achieves exact match accuracies of 18.6%, 12.2%, and 2.0% on PopQA, HotpotQA, and Musique, respectively, which are lower than the corresponding results reported in Table 3. Overall, $\text{R}^{2}$C consistently achieves the highest average performance by a substantial margin, demonstrating strong generalization across different LLM backbones.
7. Conclusions and Future Work
This paper introduces a novel and theoretically grounded UQ method for retrieval-augmented reasoning (RAR) systems, called Retrieval-Augmented Reasoning Consistency ($\text{R}^{2}$C). We argue that an effective UQ method should account for different sources of uncertainty and accurately reflect them in its final score. $\text{R}^{2}$C models uncertainty stemming from both the retriever and the generator by perturbing the reasoning process through a series of actions, including query paraphrasing, critical rethinking, and answer validation. Comprehensive experiments conducted on three datasets and five RAR models demonstrate that $\text{R}^{2}$C improves AUROC by more than 5% on average compared to state-of-the-art UQ baselines. Moreover, when used as an external signal in two downstream tasks, $\text{R}^{2}$C consistently proves effective: in Abstention, it yields around 5% gains in both F1Abstain and AccAbstain; in Model Selection, it increases exact match by approximately 7% over individual models and about 3% over selection methods. While this paper focuses on UQ for RAR models, the underlying concept of modeling and stimulating multiple sources of uncertainty is broadly applicable. Future work can extend this approach to other domains involving multiple sources of uncertainty, such as vision-language models. Moreover, in this work, we focus on short-form QA, where the final answer is an entity. Future research can explore UQ for long-form generation, which represents a more realistic scenario.
Acknowledgements.
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).
Appendix A Prompt Details
An important component of the $\text{R}^{2}$C method is the Perturbation Actions discussed in Section 4.2. These actions are implemented by prompting the LLM with specific instructions. Figures 7, 8, and 9 show the prompts used for Query Paraphrasing, Critical Rethinking, and Answer Validation, respectively. Moreover, the Answer Validation action includes a Reasoning Path Summarization component, which is also implemented through prompting an LLM. The prompt for this component is shown in Figure 10. Finally, Figure 11 presents the semantic equivalence prompt, which is used to assess the equality of two responses for majority voting and clustering in the Model Selection process described in Section 5.3, following pervious work (Bakman et al., 2024; Yaldiz et al., 2025a; Li et al., 2025a).
Figure 7. Prompt for query paraphrasing action.
Figure 8. Prompt for critical rethinking action.
Figure 9. Prompt for answer validation action.
Figure 10. Prompt for reasoning path summarization used in the answer validation action.
Figure 11. Prompt designed to evaluate the semantic equivalence between two responses to a user query.
Table 5. Abstention performance measured by the threshold-free metric AUARC. For each column, the best and second-best methods are indicated in bold and underlined, respectively. The superscript denotes a statistically significant difference compared to ReaC based on the bootstrap test ($p$ ¡ 0.05).
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) ReSearch (Chen et al., 2025a) Search-R1 (Jin et al., 2025) Avg.
Uncer. M. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq.
RrrC (Li et al., 2025a) 52.68 45.25 19.42 41.77 40.25 18.71 44.96 41.58 16.43 45.26 52.13 24.44 56.50 54.70 22.19 38.42
P(true) (Kadavath and others, 2022) 56.26 49.53 18.35 56.53 38.82 15.63 54.47 39.75 15.37 60.42 55.03 29.54 61.11 58.72 24.55 42.27
SelfC (Wang et al., 2023) 55.28 44.63 15.35 59.88 45.13 20.88 56.09 41.83 21.10 62.36 54.38 29.55 58.31 51.53 24.76 42.67
ReaC (Qi et al., 2025) 56.47 48.32 18.72 60.81 45.48 20.52 56.98 42.55 21.61 63.13 55.21 29.63 54.02 55.73 27.25 43.83
$\text{R}^{2}$C (our) 58.87 53.81 19.60 62.74 47.62 24.45 60.72 46.65 21.84 61.19 59.25 31.91 66.06 60.90 31.58 47.15
Appendix B Abstention Task
B.1. Evaluation Metrics
We evaluate $\text{R}^{2}$C on the abstention task in Section 5.2. Following Feng et al. (2024), we adopt their definitions for the evaluation metrics. Assume we have a confusion matrix with four elements, each denoted by a character: (A) Answered Correct, (B) Abstained Correct, (C) Answered Incorrect, and (D) Abstained Incorrect. Based on these, four metrics are defined for the abstention task:
  1. (1)
    Reliable Accuracy: $\frac{A}{A+C}$, measures how trustworthy the LLM’s generated (non-abstained) answers are; that is, among all answered questions, how many are correct?
  2. (2)
    Effective Reliability: $\frac{A-C}{A+B+C+D}$, balances reliability and coverage; that is, across all questions, what proportion are answered correctly minus those answered incorrectly?
  3. (3)
    Abstain Accuracy: $\frac{A+D}{A+B+C+D}$, evaluates whether abstention decisions are correct; ideally, LLMs should abstain when it would provide an incorrect answer and vice versa.
  4. (4)
    Abstain F1: the harmonic mean of precision and recall, where precision $\frac{D}{B+D}$ and recall $\frac{D}{C+D}$, providing a balanced measure between reliability and answer coverage.
In this paper, we report Abstain Accuracy and Abstain F1, as our primary goal is to evaluate the abstention capability of the uncertainty scores.
Figure 12. A case study of a successful sample from the HotpotQA dataset using the RAR model (Search-R1). The predicted answer, “Thebarton Oval,” is incorrect, and the model assigns it a relatively high uncertainty score of 0.9, indicating a desirable correlation between uncertainty and correctness.
Figure 13. A case study of a failed sample from the HotpotQA dataset using the RAR model (Search-R1). Although the predicted answer, ”Adrian Lyne,” is correct, the model assigned a relatively high uncertainty score of 0.6, revealing a mismatch between uncertainty and correctness.
B.2. Threshold Calibration
To determine the threshold $\tau_{\text{abs}}$, we perform a parameter sweep using validation sets. To construct a validation set for each dataset, we subsample 100 examples from the training set of each dataset. The only exception is PopQA, which does not have a training set. For PopQA, we instead subsample from the original test set while ensuring that our validation and test sets do not overlap. (As a reminder, as described in Section 5.1, we sample 500 examples for the test set, and the PopQA dataset consists entirely of 14K test samples.) We then generate uncertainty scores for the validation set using $\text{R}^{2}$C as well as all baseline methods.
After obtaining the validation sets and corresponding uncertainty scores, we sweep the threshold values from $0.4$ to $0.95$ with an interval of $0.05$, evaluating both AbstainAccuracy and AbstainF1. We conduct this procedure across all datasets and for all baseline methods, including RrrC, ReaC, SelfC, P(true), and our proposed approach. We first observe that both metrics exhibit similar behavioral patterns. Our results further show that all methods achieve their best performance at a threshold of $0.9$; therefore, we set $\tau_{\text{abs}}=0.9$.
B.3. Evaluation with AUARC
In Table 2, we present the performance of the abstention task using two threshold-based metrics: AbstainAccuracy and AbstainF1. The decision thresholds are determined through the detailed experiments described in Appendix B.2. However, some studies in the literature (Ren et al., 2023; Duan et al., 2025; Farquhar et al., 2024) adopt the threshold-free Area Under the Accuracy–Rejection Curve (AUARC) (Nadeem et al., 2010) as the evaluation metric. While AUARC has the advantage of being independent of a specific threshold, its final score is correlated with the model’s overall accuracy. Considering these pros and cons, we also report the abstention task results using AUARC in Table 5. We observe that the model rankings remain consistent with those in Table 2, indicating that the thresholds for AbstainAccuracy and AbstainF1 were appropriately selected. Moreover, even with the threshold-free metric, the $\text{R}^{2}$C method continues to outperform the other baselines.
Appendix C Case Study
Figure 1 illustrates the workflow of the UQ for computing the uncertainty score, which consists of two main steps: Most-likely Generation and Multi-Generations. Based on this workflow, we present two case studies—one successful and one failed, in Figures 12 and 13, respectively. In the successful case, the most-likely response is “Thebarton Oval”, which is incorrect. However, in the Multi-Generations step, only one out of ten generated responses is “Thebarton Oval”. According to Equations (LABEL:eq:mv) and (1), this results in an uncertainty score of 0.9, indicating high uncertainty. In contrast, in the failed case, the most-likely response is “Adrian Lyne”, which is correct. Yet, in the Multi-Generations step, only four out of ten responses are “Adrian Lyne”, again leading to a relatively high uncertainty score. In this sample, we observe that the answer validation action performs correctly and generates a response similar to the most-likely one. However, the query paraphrasing action produces different responses, and two of the three critical rethinking actions also yields a different result. This observation supports the discussion in Section 6.5, which highlights the potential impact of action selection configurations—whether applied per sample, per model, or per dataset.